OpenStudy (anonymous):

Will medal, Help fast please running out of time fast! Variable y varies directly with x2, and y = 32 when x = 4. Which graph represents the quadratic variation?

OpenStudy (anonymous):

This is a practice test that has no effect on the real thing, so its fine if you show me answers as long as you show how you got it.

OpenStudy (anonymous):

32=k(4)^2 k=2 A

OpenStudy (anonymous):

normally y=(1)^2, y=1 but in A (1,2) because k=2

OpenStudy (anonymous):

Wrong.